OpenStudy (anonymous):

What is the slope of the line that contains the points (2, 6) and (-1, -6)? 12 - 1 4 0 4

OpenStudy (anonymous):

\[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] you have your points, just plug and chug.

OpenStudy (anonymous):

I am not smart... This is really hard.....

OpenStudy (anonymous):

it would be 4

OpenStudy (anonymous):

...

OpenStudy (anonymous):

But how do you get the answer... Thats what makes me so confused..

OpenStudy (anonymous):

-6-6=-12 ---------- -1-2=-3 which simplifies to 4

OpenStudy (anonymous):

Your points given are (x1,y1) and (x2,y2)

OpenStudy (anonymous):

\[m=\frac{ -6-6 }{ -1-2 }\]

OpenStudy (anonymous):

Okay, thank you @tylermcmullen23 and @iambatman I kinda get it now.....

OpenStudy (anonymous):

ok good, i can help again if you need it, and im sure he can as well